how do you complete the trinomial sqaure when you are given the problem x^2-4x=8?
Answer by checkley77(12844)

x^2-4x=8
x^2-4x+4=8+4
(x-2)^2=12
x-2=sqrt12
x-2=3.4641
x=3.4641+2
x=5.4641 ans.